A decrease drop in the price of a product from R25,00 to R20,00 causes the quantity
demanded to increase from 1 500 to 2 000 units. Using the arc elasticity of demand, the
price elasticity of demand is … .
1. 0.77
2. – 0.78
3. – 0.82
4. – 1.29
1
Expert's answer
2019-02-20T14:35:07-0500

(4) – 1.29

E_d^p=((Q2-Q1)/((Q2+Q1)/2))/((P2-P1)/((P2+P1)/2))=((2000-1500)/((2000+1500)/2))/((20-25)/((20+25)/2))=0.29/(-0.22)=-1.29